OpenStudy (anonymous):

Using complete sentences, explain how you would factor completely x^9 –27. Please Help!!

OpenStudy (turingtest):

this is of the "difference of cubes" form:\[a^3-b^3=(a-b)(a^2+ab+b^2)\]so we need to identify a and b in this case, then put those numbers into this form. To do this take the cube root of each term:\[\sqrt[3]{x^9}=x^3\]\[\sqrt[3]{27}=3\]so a=x^3 and b=3 Putting this into the form above yields\[(x^3-3)(x^6+3x^3+9)\]

OpenStudy (jamesj):

Now ... by the same principle you can factor the (x^3 - 3) term

OpenStudy (anonymous):

So i would have to do that again?

OpenStudy (turingtest):

Yeah, I was thinking about that James, but do they really think an expression containing the cubed root of 3 more simplified?\[(x-\sqrt[3]{3})(x^2+x \sqrt[3]{3}+\sqrt[3]{9})(x^6+3x^3+9)\]would be how it turns out. Ugh! ugly in my opinion.

OpenStudy (jamesj):

It's definitely more 'simple', but if you want a factorization in terms of prime polynominals over the reals (which is what is being asked, I think), what you've written is it.

OpenStudy (anonymous):

Thanks guys